OpenStudy (anonymous):

n(A)=13 n(A (intersects) B)=8 n(A U B)=20 Find n(B)=

OpenStudy (zarkon):

\[n(A\cup B)=n(A)+n(B)-n(A\cap B)\]

OpenStudy (zarkon):

plug your numbers in and solve

OpenStudy (anonymous):

I don't have B though what do I do?

OpenStudy (zarkon):

\[n(A\cup B)=n(A)+n(B)-n(A\cap B)\] \[20=13+n(B)-8\] solve

OpenStudy (anonymous):

so 15 is the answer?

OpenStudy (adi3):

I think the answer is 4. not sure 20 = 13 + n(b) - 8 20 = 13 - 8 n(b) 20 = 5 + n(b) 20/5 = n(b)

OpenStudy (adi3):

@kpcoco

OpenStudy (zarkon):

15 is correct

OpenStudy (adi3):

@zarkon how did u do that.

OpenStudy (zarkon):

solve for N(B) 20=13+n(B)-8 20-13+8=n(B) 15=n(B)